# ConFIG: Towards Conflict-free Training of Physics Informed Neural Networks

### v1

@@ 4.1 Physics Informed Neural Networks
Here, $\Omega\subset R^{d}$ is a spatial domain with $\partial\Omega$ and $\overline{\Omega}$ denotes its boundary and closure, respectively. $N$ and $B$ are spatial-temporal differential operators, and $f$, $g$ and $h$ are source functions. To solve this initial boundary value problem, PINNs introduce a neural network $\hat{u}(\bm{x},t,\bm{\theta})$ for the target function $u(\bm{x},t)$, where $\theta$ is the weights of the neural networks. Then, spatial and temporal differential operators can be calculated efficiently with auto-differentiation of $N[\hat{u}(\bm{x},t,\bm{\theta}),\bm{x},t]$ and $B[\hat{u}(\bm{x},t,\bm{\theta}),\bm{x},t]$. Solving Eq. 6-8 turns to the training of the neural network with a loss function of
@@
Fig. 4 compares the performance of PINNs trained with our ConFIG method and other existing approaches. Specifically, we compare to PCGrad [16] and the IMTL-G method [27] from MTL studies as well as the LRA method [6], MinMax [11], and ReLoBRaLo [19] as established weighting methods for PINNs. The accuracy metric is the MSE between the predictions and the ground truth value on the new data points sampled in the computational domain that differ from the training data points. As the central question for all these methods is how much improvement they yield in comparison to the standard training configuration with Adam, we show the relative improvement (in percent) over the Adam baseline. Thus, +50 means half, -100 twice the error of Adam, respectively. (Absolute metrics are given in the Appendix A.9)
@@ 4.2 Multi-task Learning
While PINNs represent a highly challenging and relevant case, we also evaluate our ConFIG method for traditional Multi-Task Learning (MTL) as an outlook. We employ the widely studied CelebA dataset [32], comprising 200,000 face images annotated with 40 facial binary attributes, making it suitable for MTL with $m=40$ loss terms. Unlike PINNs, where each sub-task (loss term) carries distinct significance, and the final test loss may not reflect individual subtask performance, the CelebA experiment allows us to use the same metric to evaluate the performance for all tasks.
@@
We compared the performance of our ConFIG method with ten popular MTL baselines. Besides PCGrad and IMTL-G from before, we also compare to Linear scalarization baseline (LS) [33], Uncertainty Weighting (UW) [34], Dynamic Weight Average (DWA) [35], Gradient Sign Dropout (GradDrop) [36], Conflict-Averse Gradient Descent (CAGrad) [37], Random Loss Weighting (RLW) [38], Nash bargaining solution for MTL (Nash-MTL) [39], and Fast Adaptive Multitask Optimization (FAMO) [40]. We use two metrics to measure the performance of different methods: the mean rank metric ($MR$, lower ranks being better) [39, 40], represents the average rank of across all tasks. An $MR$ of 1 means consistently outperforming all others. In addition, we consider the average F1 score ($\overline{F_{1}}$, larger is better) to measure the average performance on all tasks.
@@
Fig. 11 presents a partial summary of the results. Our ConFIG method or M-ConFIG method emerges as the top-performing method for both $MR$ and $\overline{F_{1}}$ metric. Unlike the PINN study, we update 30 momentum variables rather than a single momentum variable for the M-ConFIG method, as indicated by the subscript ‘30.’ This adjustment is necessary because a single update step is insufficient to achieve satisfactory results in the challenging MTL experiments involving 40 tasks. Fig. 12 illustrates how the performance of the M-ConFIG method varies with different numbers of tasks and momentum-update steps in the CelebA MTL experiments. The performance of M-ConFIG tends to degrade as the number of tasks increases. This decline is attributed to longer intervals between momentum updates for each gradient, making it difficult for the momentum to accurately capture changes in the gradient. When the number of tasks becomes sufficiently large, this performance loss outweighs the benefits of M-ConFIG’s faster training speed. In our experiments, when the number of tasks reaches 10, the $\overline{F}_{1}$ score of the ConFIG method (0.635) already surpasses that of M-ConFIG (0.536) with the same training time. However, increasing the number of iterations can effectively mitigate this performance degradation. In the standard 40-task MTL scenario, increasing the number of gradient update steps per iteration to 20 allows M-ConFIG to achieve a performance level comparable to ConFIG while reducing the training time to only 56% of that required by ConFIG. Notably, alternating momentum updates can sometimes even outperform the ConFIG method with the same number of training epochs, as seen in the $n_{updates}=30$ case in the MTL experiment, as well as in the Burgers and Kovasznay cases in the two-loss PINNs experiment.
@@
A potential limitation of the proposed method is the performance degradation of the M-ConFIG method as the number of losses increases. Additionally, the memory consumption required to store the momentum for each gradient grows with both the size of the parameter space and the number of loss terms. As discussed in the MTL experiment, one approach to mitigate the performance degradation is to update not just a single gradient but a stochastic subset of gradients during each iteration, albeit at the cost of an increased computational budget. Nonetheless, further investigation and improvement of the M-ConFIG method’s performance in scenarios with a large number of loss terms is a promising direction for future research, aiming to broaden the applicability of the ConFIG approach.
@@ 5 Conclusions
In this study, we presented ConFIG, a method designed to alleviate training conflicts between loss terms with distinct behavior. The ConFIG method achieves conflict-free training by ensuring a positive dot product between the final gradient and each loss-specific gradient. Additionally, we introduce a momentum-based approach, replacing the gradient with alternately updated momentum for highly efficient iterations. Empirically, the proposed methods have clearly shown a superior performance compared with existing training strategies for PINNs, and even outperform SOTA methods for challenging MTL scenarios.
Acknowledgments and Disclosure of Funding
@@ Appendix A Appendix
A.1 Convergence Analysis of ConFIG method
In this section, we will give a convergence analysis for both convex (Theorem 1) and non-convex landscapes (Theorem 2) for our ConFIG method.
@@ A.4 Proof of the Equivalence
To prove the equivalence between Eq. 23 and Eq. 45 is equivalent to prove that $\mathcal{U}\left[\mathcal{O}(\bm{g}_{1},\bm{g}_{2})\right]+\mathcal{U}\left[\mathcal{O}(\bm{g}_{2},\bm{g}_{1})\right]=k[\mathcal{U}(\bm{g}_{1}),\mathcal{U}(\bm{g}_{2})]^{-\top}\mathbf{1}_{m}$ where k>0. Thus, we can first calculate
@@ A.5 M-Config and Momentum Strategies
In the current study, we use the Adam optimizer to train baseline PINNs, as shown in Algorithm 3. The green line in the algorithm can be replaced by other methods like PCGrad or the IMTL-G method. The Adam optimizer uses the first momentum $m$ to average the recent update direction and the second momentum $v$ to rescale the learning rate of each parameter. If we want to use the momentum to replace the gradient for the ConFIG update, an intuitive approach could be to calculate the first and second momentum for each loss-specific gradient and use the final rescaled momentum to calculate the input for the ConFIG update, as shown in "MA-ConFIG" in Algorithm 4. However, this approach could result in several issues:
@@
Table 1: Test results of different momentum configurations. "MA-ConFIG" refers to the strategy with the second momentum inside the ConFIG operation, and "NaN" means the training failed due to numerical stability.
@@ A.6 Computational Cost
It is also worth noting that the time cost of the ConFIG method is comparable to other gradient-based methods, indicating that the pseudoinverse operation of the ConFIG method does not add significant computational overhead. This study uses the PyTorch implementation, which utilizes singular value decomposition (SVD) to calculate the pseudoinverse. Although the exact computational cost depends on the specific numerical algorithm, the general time complexity of SVD is $\mathcal{O}(nm^{2})$ [41] where $n$ is the number of gradient elements (i.e., the number of neural network’s parameter) and $m$ is the number of gradients in the current study ($n>m$).
This is favorable since the number of network parameters is usually much larger than the number of loss terms, and the time complexity of SVD scales linearly with the former.
@@ A.7 PDE Details
No analytical solution is available for Burgers equation with these given boundary conditions. In the current study, we use numerical solutions from PhiFlow [42] as the ground truth solution to evaluate the PINNs’ performance.
@@
Similarly, no analytical solution is available for the 1D Schrödinger equation with given boundary conditions. We use the numerical solution provided by deepXDE [43] to evaluate the performance of trained PINNs.
@@
where $\bm{u}$ is the velocity vector, $p$ is the pressure and $Re$ is the Reynolds number. Kovasznay flow is a steady case in two space dimensions where the transient term $\partial\bm{u}/\partial t$ equals 0. The analytical solution of the velocity $\bm{u}(x,y)=[u_{x}(x,y),u_{y}(x,y)]$ and pressure $p(x,y)$ is [44, 45]
@@
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$,$\quad\nu=\frac{1}{\operatorname{Re}}=\frac{1}{40}$, $x\in[-0.5,1]$ and $y\in[-0.5,1.5]$. Meanwhile, Beltrami flow is an unsteady case in three space dimensions where the velocity $\bm{u}(x,y,z,t)=[u_{x}(x,y,z,t),u_{y}(x,y,z,t),u_{z}(x,y,z,t)]$ and pressure $p(x,y,z,t)$ follows the analytical solution of [46]
@@ A.9 PINN Evaluations
$[\mathcal{L}_{\mathcal{N}},\mathcal{L}_{\mathcal{BI}}]$
@@
$[\mathcal{L}_{\mathcal{N}},\mathcal{L}_{\mathcal{BI}}]$
@@ A.10 Results of CelebA MTL Experiments
A.11 Training Details

### v2

@@ 4.1 Physics Informed Neural Networks
Here, $\Omega\subset R^{d}$ is a spatial domain with $\partial\Omega$ and $\overline{\Omega}$ denotes its boundary and closure, respectively. $N$ and $B$ are spatial-temporal differential operators, and $f$, $g$ and $h$ are source functions. To solve this initial boundary value problem, PINNs introduce a neural network $\hat{u}(\bm{x},t,\bm{\theta})$ for the target function $u(\bm{x},t)$, where $\theta$ is the weights of the neural networks. Then, spatial and temporal differential operators can be calculated efficiently with auto-differentiation of $N[\hat{u}(\bm{x},t,\bm{\theta}),\bm{x},t]$ and $B[\hat{u}(\bm{x},t,\bm{\theta}),\bm{x},t]$. Solving Eq. 6-8 turns to the training of neural networks with a loss function of
@@
Fig. 4 compares the performance of PINNs trained with our ConFIG method and other existing approaches. Specifically, we compare to PCGrad (Yu et al., 2020) and the IMTL-G method (Liu et al., 2021b) from MTL studies as well as the LRA method (Wang et al., 2021), MinMax (Liu & Wang, 2021), and ReLoBRaLo (Bischof & Kraus, 2021) as established weighting methods for PINNs. The accuracy metric is the MSE between the predictions and the ground truth value on the new data points sampled in the computational domain that differ from the training data points. As the central question for all these methods is how much improvement they yield in comparison to the standard training configuration with Adam, we show the relative improvement (in percent) over the Adam baseline. Thus, +50 means half, -100 twice the error of Adam, respectively. (Absolute metrics are given in the Appendix A.9)
@@
Additional Tests on Challenging PDEs
Recently, several benchmark problems have been proposed as challenging tasks for training PINNs  (Hao et al., 2023). We have also tested our ConFIG and M-ConFIG methods together with other methods on these challenging tasks. The results are summarized in Appendix. A.11. They show that, although our methods do not resolve all inherent difficulties of PINN training, our methods still consistently outperform the other methods. These results shows the general appeal of the proposed methods for challenging, and high-dimensional training scenarios.
@@ 4.2 Multi-task Learning
While PINNs represent a highly challenging and relevant case, we also evaluate our ConFIG method for traditional Multi-Task Learning (MTL) as an outlook. We employ the widely studied CelebA dataset (Liu et al., 2015), comprising 200,000 face images annotated with 40 facial binary attributes, making it suitable for MTL with $m=40$ loss terms. Unlike PINNs, where each sub-task (loss term) carries distinct significance, and the final test loss may not reflect individual subtask performance, the CelebA experiment allows us to use the same metric to evaluate the performance for all tasks.
@@
We compared the performance of our ConFIG method with ten popular MTL baselines. Besides PCGrad and IMTL-G from before, we also compare to Linear scalarization baseline (LS) (Caruana, 1997), Uncertainty Weighting (UW) (Cipolla et al., 2018), Dynamic Weight Average (DWA) (Liu et al., 2019), Gradient Sign Dropout (GradDrop) (Chen et al., 2020), Conflict-Averse Gradient Descent (CAGrad) (Liu et al., 2021a), Random Loss Weighting (RLW) (Lin et al., 2022), Nash bargaining solution for MTL (Nash-MTL) (Navon et al., 2022), and Fast Adaptive Multitask Optimization (FAMO) (Liu et al., 2023). We use two metrics to measure the performance of different methods: the mean rank metric ($MR$, lower ranks being better) (Navon et al., 2022; Liu et al., 2023), represents the average rank of across all tasks. An $MR$ of 1 means consistently outperforming all others. In addition, we consider the average F1 score ($\overline{F_{1}}$, larger is better) to measure the average performance on all tasks.
Fig. 11 presents a partial summary of the results. Our ConFIG method or M-ConFIG method emerges as the top-performing method for both $MR$ and $\overline{F_{1}}$ metric. Unlike the PINN study, we update 30 momentum variables rather than a single momentum variable for the M-ConFIG method, as indicated by the subscript ‘30.’ This adjustment is necessary because a single update step is insufficient to achieve satisfactory results in the challenging MTL experiments involving 40 tasks.
Varying Number of Tasks
@@
Fig. 12 illustrates how the performance of the M-ConFIG method varies with different numbers of tasks and momentum-update steps in the CelebA MTL experiments. The performance of M-ConFIG tends to degrade as the number of tasks increases. This decline is attributed to longer intervals between momentum updates for each gradient, making it difficult for the momentum to accurately capture changes in the gradient. When the number of tasks becomes sufficiently large, this performance loss outweighs the benefits of M-ConFIG’s faster training speed. In our experiments, when the number of tasks reaches 10, the $\overline{F}_{1}$ score of the ConFIG method (0.635) already surpasses that of M-ConFIG (0.536) with the same training time. However, increasing the number of iterations can effectively mitigate this performance degradation. In the standard 40-task MTL scenario, increasing the number of gradient update steps per iteration to 20 allows M-ConFIG to achieve a performance level comparable to ConFIG while reducing the training time to only 56% of that required by ConFIG. Notably, alternating momentum updates can sometimes even outperform the ConFIG method with the same number of training epochs, as seen in the $n_{updates}=30$ case in the MTL experiment, as well as in the Burgers and Kovasznay cases in the two-loss PINNs experiment.
@@
A potential limitation of the proposed method is the potential performance degradation of the M-ConFIG method as the number of loss terms increases. Additionally, the memory consumption required to store the momentum for each gradient grows with both the size of the parameter space and the number of loss terms. As discussed in the MTL experiment, one approach to mitigate the performance degradation is to update not just a single gradient but a stochastic subset of gradients during each iteration, albeit at the cost of an increased computational budget. Nonetheless, further investigation and improvement of the M-ConFIG method’s performance in scenarios with a large number of loss terms is a promising direction for future research, aiming to broaden the applicability of the ConFIG approach. Meanwhile, we also notice that there are still many challenges in the training of PINNs, e.g., chaotic problems, that can not be addressed by eliminating the conflicts between different loss terms during the training. Future efforts, like improved network structures and imposed causality, will be required to further improve PINN training.
@@ 5 Conclusions
In this study, we have presented ConFIG, a method designed to alleviate training conflicts between loss terms with distinct behavior. The ConFIG method achieves conflict-free training by ensuring a positive dot product between the final gradient and each loss-specific gradient. Additionally, we introduce a momentum-based approach, replacing the gradient with alternately updated momentum for highly efficient iterations. In our experiments the proposed methods have shown superior performance compared with a wide range of existing training strategies for PINNs. The config variants even outperform SOTA methods for challenging MTL scenarios.
Acknowledgments
@@ Appendix A Appendix
A.1 Convergence Proof for the ConFIG method
In this section, we will give the proof for convergence for both convex (Theorem 1) and non-convex landscapes (Theorem 2) for our ConFIG method.
@@ A.4 Proof of the Equivalence
To prove the equivalence between Eq. 23 and Eq. 45 is equivalent to prove that $\mathcal{U}\left[\mathcal{O}(\bm{g}_{1},\bm{g}_{2})\right]+\mathcal{U}\left[\mathcal{O}(\bm{g}_{2},\bm{g}_{1})\right]=k[\mathcal{U}(\bm{g}_{1}),\mathcal{U}(\bm{g}_{2})]^{-\top}\mathbf{1}_{m}$ where k¿0. Thus, we can first calculate
@@ A.5 M-Config and Momentum Strategies
In the current study, we use the Adam optimizer to train baseline PINNs, as shown in Algorithm 3. The green line in the algorithm can be replaced by other methods like PCGrad or the IMTL-G method. The Adam optimizer uses the first momentum $m$ to average the recent update direction and the second momentum $v$ to rescale the learning rate of each parameter. If we want to use the momentum to replace the gradient for the ConFIG update, an intuitive approach could be to calculate the first and second momentum for each loss-specific gradient and use the final rescaled momentum to calculate the input for the ConFIG update, as shown in ”MA-ConFIG” in Algorithm 4. However, this approach could result in several issues:
@@
Table 1: Test results of different momentum configurations. ”MA-ConFIG” refers to the strategy with the second momentum inside the ConFIG operation, and ”NaN” means the training failed due to numerical stability.
@@ A.6 Computational Cost
It is also worth noting that the time cost of the ConFIG method is comparable to other gradient-based methods, indicating that the pseudoinverse operation of the ConFIG method does not add significant computational overhead. This study uses the PyTorch implementation, which utilizes singular value decomposition (SVD) to calculate the pseudoinverse. Although the exact computational cost depends on the specific numerical algorithm, the general time complexity of SVD is $\mathcal{O}(nm^{2})$ (Grasedyck, 2010) where $n$ is the number of gradient elements (i.e., the number of neural network’s parameter) and $m$ is the number of gradients in the current study ($n>m$). This is favorable since the number of network parameters is usually much larger than the number of loss terms, and the time complexity of SVD scales linearly with the former.
On the side of memory, all gradient-based methods, including PCGrad, IMTL-G, and our ConFIG method, usually require more memory compared to the weighting method during training due to the additional cost of storing loss-specific gradients for each loss term in each iteration. The complexity of this additional memory is $\mathcal{O}(nm)$. Meanwhile, although momentum acceleration helps to increase the training speed, it also introduces an increased memory cost with a complexity of $\mathcal{O}(nm)$, as it also needs to store the first and second momentum during the training procedure. This could be a potential challenge when dealing with large-scale problems.
@@ A.7 PDE Details
No analytical solution is available for Burgers equation with these given boundary conditions. In the current study, we use numerical solutions from PhiFlow (Holl et al., 2020) as the ground truth solution to evaluate the PINNs’ performance.
@@
Similarly, no analytical solution is available for the 1D Schrödinger equation with given boundary conditions. We use the numerical solution provided by deepXDE (Bajaj et al., 2023) to evaluate the performance of trained PINNs.
@@
where $\bm{u}$ is the velocity vector, $p$ is the pressure and $Re$ is the Reynolds number. Kovasznay flow is a steady case in two space dimensions where the transient term $\partial\bm{u}/\partial t$ equals 0. The analytical solution of the velocity $\bm{u}(x,y)=[u_{x}(x,y),u_{y}(x,y)]$ and pressure $p(x,y)$ is (Kovasznay, 1948; Truesdell, 1954)
@@
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$,$\quad\nu=\frac{1}{\operatorname{Re}}=\frac{1}{40}$, $x\in[-0.5,1]$ and $y\in[-0.5,1.5]$. Meanwhile, Beltrami flow is an unsteady case in three space dimensions where the velocity $\bm{u}(x,y,z,t)=[u_{x}(x,y,z,t),u_{y}(x,y,z,t),u_{z}(x,y,z,t)]$ and pressure $p(x,y,z,t)$ follows the analytical solution of (Gromeka, 1881)
@@ A.9 PINN Evaluations
$[\mathcal{L}_{\mathcal{N}},\mathcal{L}_{\mathcal{B}}]$
@@
$[\mathcal{L}_{\mathcal{N}},\mathcal{L}_{\mathcal{B}}]$
@@ A.10 Results of CelebA MTL Experiments
A.11 Additional results on challenging PDEs
To evaluate the performance of our method for challenging and high-dimensional test cases, we apply it to three problems from a recent benchmark for PINNs  (Hao et al., 2023). In the following, we perform a comprehensive investigation of the performance of our method in solving these problems.
A.11.1 High dimensional PDEs
We introduce the N-dimensional Poisson equation as the test case for high-dimensional problems. The governing PDE is
$-\nabla^{2}u=\frac{\pi^{2}}{4}\sum_{i=1}^{n}\sin(\frac{\pi}{2}x_{i}),$ (44)
where $n$ is the number of dimensionality. In the current experiment, we choose $n=5$ and set the spatial domain as $x\in[0,1]^{5}$ following the configuration of Hao et al. (2023). The ground truth solution is
$u=\sum_{i=1}^{n}\sin(\frac{\pi}{2}x_{i}).$ (45)
The PNd problem employs Dirichlet boundary conditions with the boundary value equal to the analytical solution.
Tab. 20 and Fig. 23 summarize the test results of different methods. Our M-ConFIG method ranks first among all methods, followed by the ConFIG methods, showing the superiority of our methods in dealing with high-dimensional problems.
Table 20: Test MSE of PINNs trained for PNd equation. All values ×$\times 10^{-6}$.
$[\mathcal{L}_{\mathcal{N}},\mathcal{L}_{\mathcal{B}}]$
Adam $1.916\pm 0.284$
PCGrad $1.313\pm 0.097$
IMTL-G $0.520\pm 0.123$
MinMax $4.604\pm 0.271$
ReLoBRaLo $2.265\pm 0.263$
LRA $0.639\pm 0.340$
ConFIG $0.461\pm 0.141$
M-ConFIG $\mathbf{0.415\pm 0.052}$
Figure 23: The solution domain of PNd problem on $[x,y,0.5,0.5,0.5]$ plane.
A.11.2 Multi-scale Problems
We choose the multi-scale heat transfer (HeatMS) problem as a strongly anisotropic test case. We set different heat-transfer coefficients in different spatial directions to give the solution different scales in each direction. Following the configuration of Hao et al. (2023), the governing equation is
$\frac{\partial u}{\partial t}-\frac{1}{(500\pi)^{2}}\frac{\partial^{2}u}{\partial x^{2}}-\frac{1}{\pi^{2}}\frac{\partial^{2}u}{\partial y^{2}}=0,$ (46)
with the initial condition of $u(x,y,0)=\sin(20\pi x)\sin(\pi y)$ and boundary condition of $u(x,y,t)=0$. The spatial time domain is $x\in[0,1]$, $y\in[0,1]$ and $t\in[0,5]$. Fig. 24 illustrates a sample solution of the ground truth when $t=3.0s$.
Tab. 21 and Fig. 25 show the predictions by different methods. While all of the methods struggle to fully resolve the anisotropic solutions, as shown in Fig. 25, our M-ConFIG method still ranks first. This test case illustrates that the scaling issues of anisotropic PDEs can not be solved purely by finding a better balance between different loss terms. Nonetheless, ConFIG fares on-par with other methods, and the momentum terms of M-ConFIG help to partially address the scaling of the HeatMS test case.
Table 21: Test MSE of PINNs trained for HeatMS problem. All values are scaled with $10^{-3}$.
Adam $7.612\pm 0.004$
PCGrad $7.600\pm 0.012$ $7.603\pm 0.012$
IMTL-G $6.693\pm 0.241$ $38.529\pm 43.666$
MinMax $7.595\pm 0.055$ $7.243\pm 0.458$
ReLoBRaLo $7.585\pm 0.039$ $7.558\pm 0.041$
LRA $7.654\pm 0.004$ $7.652\pm 0.000$
ConFIG $7.529\pm 0.074$ $7.585\pm 0.020$
M-ConFIG $\mathbf{5.978\pm 0.092}$ $\mathbf{7.147\pm 0.001}$
Figure 24: The ground truth solution of HeatMS problem when $t=3.0s$.
Figure 25: The predictions of PINNs for HeatMS problem when $t=3.0s$ (two-loss scenario, values scaled with $10^{-2}$).
A.11.3 Chaotic Problems
To test the potential limitations of the proposed methods, we target the KS equation as a representative of chaotic problems. Starting from an analytic initial state, the solution of the KS equation will gradually develop into a chaotic state due to the strong nonlinearity, dissipative, and destabilizing effects. Following the configuration of Hao et al. (2023), the governing equation is
$\frac{\partial u}{\partial t}+\frac{100}{16}u\frac{\partial u}{\partial x}+\frac{100}{16^{2}}\frac{\partial^{2}u}{\partial x^{2}}+\frac{100}{16^{4}}\frac{\partial^{4}u}{\partial x^{4}}=0,$ (47)
with an initial condition of $u(x,0)=\cos(x)(1+\sin(x))$ and periodic boundary conditions. The spatial time domain is $x\in[0,2\pi]$ and $t\in[0,1]$.
Table 22: Test MSE of PINNs for the KS equation.
Adam $1.094\pm 0.002$ $1.094\pm 0.002$
PCGrad $1.089\pm 0.008$ $1.061\pm 0.002$
IMTL-G $1.110\pm 0.016$ $1.089\pm 0.031$
MinMax $1.072\pm 0.000$ $1.084\pm 0.007$
ReLoBRaLo $1.074\pm 0.019$ $1.080\pm 0.016$
LRA $1.099\pm 0.039$ $1.098\pm 0.027$
ConFIG $1.089\pm 0.004$ $1.062\pm 0.007$
M-ConFIG $1.123\pm 0.011$ $1.118\pm 0.008$
Tab. 22 shows the predictions by different methods. The evaluation shows that no methods succeeds to capture the transition of chaos and generate smooth solutions. This agrees with the conclusions in Hao et al. (2023)’s, where likewise no method under consideration manages to converge to an accurate solution.
The inherent challenge of solving the KS equations is a combination of several factors. On the one hand, the negative diffusion term $\partial^{2}u/\partial x^{2}$ injects energy into the system, inducing instability and complex structures. On the other hand, The dissipative term $\partial^{4}u/\partial x^{4}$ removes energy at small scales, formulating energy cascade across different scales. This energy cascade causes huge difficulties for numerical methods since the fine-scale structures must be resolved accurately. From the PINN side, this requires the network to have the ability to estimate the derivatives accurately on different scales via auto differentiation. The balance between different loss terms, as all the benchmark methods are trying to achieve, is not helpful in addressing this fundamental issue. Hence, this test case serves as a failure case, highlighting that improved optimizers alone do not suffice to address these challenges. It will be an interesting topic of future work to evaluate their effectiveness in combination with other changes, such as improvements on the architecture side.
A.12 Training Details